# Graph-Informed Adversarial Modeling: Infimal Subadditivity of Interpolative Divergences

### 20 Mar 2026 v1

@@ 5.1 Infimal subadditivity for Bayesian networks over a DAG
$\begin{split}I_{1,2}[\phi](x_{1},x_{2})+I_{1,2}[\psi](x_{1}^{\prime},x_{2}^{\prime})&=\int_{\mathcal{X}_{3}}\phi(x_{1},x_{2},x_{3})Q_{3\mid 2}(dx_{3})+\int_{\mathcal{X}_{3}}\psi(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime})Q_{3\mid 2}(dx_{3}^{\prime})\\ &=\int_{\mathcal{X}_{3}}\int_{\mathcal{X}_{3}}(\phi(x_{1},x_{2},x_{3})+\psi(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}))Q_{3\mid 2}(dx_{3})Q_{3\mid 2}(dx_{3}^{\prime})\\ &\leq\int_{\mathcal{X}_{3}}\int_{\mathcal{X}_{3}}c(\bm{x},\bm{x}^{\prime})Q_{3\mid 2}(dx_{3})Q_{3\mid 2}(dx_{3}^{\prime})\end{split}$
Then, the result follows as in proof of Theorem 5.4. The proof of the second part of the theorem is based on Theorem 3.7 in Baptista et al. (2025). Specifically,
$\begin{split}\mathfrak{D}^{p}_{\rm{KL},\varepsilon}(Q\|P)&=\mathfrak{D}^{p}_{\rm{KL},\varepsilon}(Q_{\{1,2\}}\otimes Q_{3\mid 2}\|P_{\{1,2\}}\otimes Q_{3\mid 2})\\ &\leq\inf_{R_{\{1,2\}}\otimes R_{3\mid 2}}\inf_{\begin{subarray}{c}\pi\in\Pi(Q_{\{1,2\}},R_{\{1,2\}}),\\ \tilde{\pi}\in\Pi(Q_{3\mid 2},R_{3\mid 2})\end{subarray}}\Bigg\{\int\|\bm{x}_{\{1,2\}}-\bm{y}_{\{1,2\}}\|_{p}^{p}d\pi+\varepsilon\mathfrak{D}_{\rm{KL}}(R_{\{1,2\}}\|P_{\{1,2\}})\\ &\hskip 85.35826pt+\int\|x_{3}-y_{3}\|_{p}^{p}d\tilde{\pi}+\varepsilon\mathfrak{D}_{\rm{KL}}(R_{3\mid 2}\|Q_{3\mid 2})\Bigg\}\\ &\leq\mathfrak{D}^{p}_{\rm{KL},\varepsilon}(Q_{\{1,2\}}\|P_{\{1,2\}})+\mathfrak{D}^{p}_{\rm{KL},\varepsilon}(Q_{3\mid 2}\|Q_{3\mid 2})=\mathfrak{D}^{p}_{\rm{KL},\varepsilon}(Q_{\{1,2\}}\|P_{\{1,2\}})\end{split}$
For the other direction, we work exactly the same as in the proof of Theorem 3.7 in Baptista et al. (2025). Based on the above, one can generalize to any DAG, similar to the proof of Theorem 5.4. ∎
From the perspective of adversarial learning, the preceding theorems justify replacing a graph-agnostic GAN with a monolithic discriminator by a graph-informed GAN with family-level discriminators. This is not only a computational simplification; since GAN training amounts to a learned two-sample testing problem (Goodfellow et al., 2014; Nowozin et al., 2016), localization decomposes a high-dimensional discrimination task into a collection of lower-dimensional subproblems that are easier to optimize and diagnose. Theorems 5.4, 5.6, and 5.7 show that these local games define principled surrogates for the corresponding graph-constrained global objectives.
@@
The same lower and upper bounds hold for IPMs. This follows directly from the monotonicity property $\mathfrak{D}_{f}^{\Gamma_{i\cup\mathrm{Pa}(i)}}(Q_{i\cup\mathrm{Pa}(i)}\|P_{i\cup\mathrm{Pa}(i)})\leq\mathfrak{D}_{f}^{\Gamma}(Q\|P)$ for all $i\in V$. The proof of the upper bound is already contained in the proof of Theorem 5.4. Next, we state a lemma that provides sufficient conditions for subadditivity of a function and will be used in Remark 5.9 and Example 5.10.
Lemma 5.8.
If $F:[0,\infty)\to\mathbb{R}$ is concave and $F(0)=0$ (or more general $F(0)\geq 0$), then $F$ satisfies sub-additivity, i.e $F(x+y)\leq F(x)+F(y)$ for every $x,y\in[0,\infty)$.
Proof.
It is enough to observe that
$F(x)$ $=F(\frac{y}{x+y}0+\frac{x}{x+y}(x+y))\geq\frac{x}{x+y}F(x+y)$
$F(y)$ $=F(\frac{x}{x+y}0+\frac{y}{x+y}(x+y))\geq\frac{y}{x+y}F(x+y)$
By summing the above inequalities, we obtain sub-additivity, i.e.,
$F(x)+F(y)\geq\frac{x}{x+y}F(x+y)+\frac{y}{x+y}F(x+y)=F(x+y)$
@@
Furthermore, since $f^{*}$ in the variational representation of the $(f,\Gamma)$-divergences is concave and satisfies $f^{*}(0)\geq 0$ (see Lemma 42 in Birrell et al. (2022a)), by Lemma 5.8, the following subadditivity property for $(f,\Gamma)$-divergences holds
$\mathfrak{D}_{f}^{\oplus_{i=1}^{n}\Gamma_{i\,\cup\,\mathrm{Pa}(i)}}\left(Q_{i\,\cup\,\mathrm{Pa}(i)}\middle\|P_{i\,\cup\,\mathrm{Pa}(i)}\right)\leq\sum_{i=1}^{n}\mathfrak{D}_{f}^{\Gamma_{i\,\cup\,\mathrm{Pa}(i)}}\left(Q_{i\,\cup\,\mathrm{Pa}(i)}\middle\|P_{i\,\cup\,\mathrm{Pa}(i)}\right).$
(i.e., using the fact that $f^{*}(\sum_{i=1}^{n}\gamma_{i}(x_{i},\bm{x}_{\mathrm{Pa}(i)}))\leq-\sum_{i=1}^{n}f^{*}(\gamma_{i}(x_{i},\bm{x}_{\mathrm{Pa}(i)}))$).
Example 5.10 (Subadditivity for KL-divergence via variational representation).
In Section 4.1 of Ding et al. (2021), the authors establish subadditivity (see Definition 5.1) of the KL divergence on Bayesian networks over an arbitrary DAG $G$, by relying on the definition equation 3. In contrast, in this example, we re-establish the subadditivity of the KL divergence on Bayesian networks using its variational characterization.
Precisely, let $P,Q\in\mathcal{P}^{G}$. Under appropriate assumptions (see Broniatowski and Keziou (2006)), the solution of equation 4 for $f=f_{\mathrm{KL}}$, i.e, $\mathfrak{D}_{\rm{KL}}(Q\|P)=\sup_{\gamma\in\mathcal{M}_{b}(\mathcal{X})}\left\{E_{Q}[\gamma]-E_{P}[e^{\gamma-1}]\right\}$ is given by $\gamma^{*}=\log\left(\frac{dQ}{dP}\right)$. By the factorization of $P$ and $Q$ and it is straightforward to check that $\gamma^{*}\in\bar{\Gamma}=\oplus_{i=1}^{n}\mathcal{M}_{b}(\mathcal{X}_{i\,\cup\,\mathrm{Pa}(i)})$ because
$\gamma^{*}=\sum_{i=1}^{n}\log\frac{Q_{i\mid\mathrm{Pa}(i)}}{P_{i\mid\mathrm{Pa}(i)}},\qquad Q\text{-a.s.}$ (31)
Therefore, $\mathfrak{D}_{\rm{KL}}(Q\|P)=\mathfrak{D}_{\rm{KL}}^{\bar{\Gamma}}(Q\|P)$. We use Lemma 5.8 for $f^{*}_{\rm{KL}}$ and bound as follows:
$\begin{split}\mathfrak{D}_{\rm{KL}}^{\bar{\Gamma}}(Q\|P)&\leq\sup_{\begin{subarray}{c}\bar{\gamma}=\sum_{i=1}^{n}h_{i}(x_{i},x_{\mathrm{Pa}(i)}),\\ h_{i}\in\mathcal{M}_{b}(\mathcal{X}_{i\,\cup\,\mathrm{Pa}(i)})\end{subarray}}\left\{\sum_{i=1}^{n}E_{Q}[h_{i}]-E_{P}[e^{\sum_{i=1}^{n}h_{i}-1}]\right\}\\ &\leq\sum_{i=1}^{n}\sup_{h_{i}\in\mathcal{M}_{b}(\mathcal{X}_{i\,\cup\,\mathrm{Pa}(i)})}\left\{E_{Q_{i\,\cup\,\mathrm{Pa}(i)}}[h_{i}]-E_{P_{i\,\cup\,\mathrm{Pa}(i)}}[e^{h_{i}-1}]\right\}\\ &\leq\sum_{i=1}^{n}\mathfrak{D}_{\textrm{KL}}(Q_{i\,\cup\,\mathrm{Pa}(i)}\|P_{i\,\cup\,\mathrm{Pa}(i)})\end{split}$
Remark 5.11 (We cannot have subadditivity in general).
The optimizer $\gamma^{*}$ does not, in general, belong to the subspace $\bar{\Gamma}\subset\Gamma$ which would be required to invoke subadditivity via concavity. As a result, one cannot directly apply Lemma 5.8, which leverages the concavity of $f^{*}$ , to establish the desired bound. In fact, let $\bm{X}=(X_{1},X_{2},X_{3})$ and $R^{*}$ denote the optimal intermediate measure that solves equation 7. Indeed, by the properties of $(f,\Gamma)$-divergences and equation 8,
$R^{*}\propto(f_{\rm{KL}}^{*})^{\prime}(\gamma^{*})dP=e^{\gamma^{*}-1}dP$
which together with the factorization of $Q$, we obtain
By explicitly computing the conditional distribution $R^{*}_{3\mid\mathrm{Pa}^{R^{*}}\!\!(3)}(dx_{3})$, where $\mathrm{Pa}^{R^{*}}\!\!(3)$ denotes the set of parents of node 3 under the distribution $R^{*}$, we obtain
$R^{*}_{3\mid\mathrm{Pa}^{R^{*}}\!\!(3)}(dx_{3})=\frac{e^{\gamma^{*}(x_{1},x_{2},x_{3})-1}\,P_{3\mid 2}(dx_{3})}{E_{P_{3\mid 2}}\left[e^{\gamma^{*}-1}\right]}.$
This expression reveals that $\mathrm{Pa}^{R^{*}}\!\!(3)=\{1,2\}\neq\{2\}=\mathrm{Pa}^{Q}(3)$. Consequently, the optimal intermediate measure $R^{*}$ does not share the same conditional independence structure as $Q$, which prevents the likelihood terms from being paired in the same manner as in equation 31.
@@ 5.2 Examples of admissible sets satisfying condition (C)
Let $G$ be an arbitrary DAG. Then, $f$-divergences satisfy $\frac{1}{n}$-linear generalized infimal additivity
@@
for $\bm{x}=(x_{1},\dots,x_{n})$ and $\bm{x}^{\prime}=(x^{\prime}_{1},\dots,x^{\prime}_{n})$. Let $\Gamma^{L}=\textrm{Lip}^{L}_{b}(\mathcal{X})$. In this case, condition (C) can be satisfied if $Q_{V^{(i)}\mid\mathrm{Pa}(i)}(d{\bm{x}}_{V^{(i)}})$ with $V^{(i)}=V\setminus[i\cup\mathrm{Pa}(i)]$ viewed as a function of $\bm{x}_{\mathrm{Pa}(i)}$, is Lipschitz-continuous with respect to 1-Wasserstein distance (Bartl and Eckstein, 2025; Piening and Chung, 2024; Altekrüger et al., 2023; Bénézet et al., 2024). We prove this result in Proposition B.1.
In generative modeling, the assumptions given in equation 39 can be interpreted as follows: Let $(\mathcal{Z},\mathcal{F},\nu)$ be a probability space, $\mathcal{X}_{\mathrm{Pa}(i)}\subset\mathbb{R}^{m}$ and $g:\mathcal{X}_{\mathrm{Pa}(i)}\times\mathcal{Z}\to\mathbb{R}^{d}$. We assume that for every $\bm{x}_{\mathrm{Pa}(i)}\in\mathcal{X}_{\mathrm{Pa}(i)}$, the map $z\mapsto g(\bm{x}_{\mathrm{Pa}(i)},z)$ is measurable and $g(\bm{x}_{\mathrm{Pa}(i)},\cdot)_{\#}\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$, and for any $\bm{x}_{\mathrm{Pa}(i)},\bm{x}^{\prime}_{\mathrm{Pa}(i)}\in\mathcal{X}_{\mathrm{Pa}(i)}$ and all $z\in\mathcal{Z}$, $\|g(\bm{x}_{\mathrm{Pa}(i)},z)-g(\bm{x}^{\prime}_{\mathrm{Pa}(i)},z)\|\leq Cd_{\mathcal{X}_{\mathrm{Pa}(i)}}(\bm{x}_{\mathrm{Pa}(i)},\bm{x}^{\prime}_{\mathrm{Pa}(i)})$. Then, a measurable generator with finite first moment and uniform Lipschitz dependence on the conditioning random vector induces a $W_{1}$-Lipschitz family of conditional distributions, see Arjovsky et al. (2017); Papamakarios et al. (2021). This is proved in Appendix B.
Therefore, the assumption given equation 39 expresses a quantitative stability property of the model: small perturbations of the conditioning variables induce proportionally small changes in the corresponding conditional distributions in $1$-Wasserstein distance. This stability naturally arises when conditional distributions are represented as the pushforwards of a reference noise distribution through generators that are Lipschitz in their conditioning variables.
@@ 6 Infimal subadditivity in action
Section 5 shows that when the target distribution factorizes according to a known Bayesian network over a graph $G$, the corresponding global objective function can be related to a sum of lower-dimensional objective functionals defined on graph-informed families. In adversarial learning terms, this suggests replacing a graph-agnostic GAN with a monolithic discriminator by a graph-informed GAN with localized discriminators acting on families of variables. The aim of this section is to show how the graph-informed strategy behaves in practice across a range of structured target distributions. In particular, we find that graph-informed GANs often improve optimization stability and the recovery of the underlying conditional structure, while also revealing trade-offs between family-level and global distributional discrepancy, and between predictive accuracy and computational cost.
@@ 22   Record validation and runtime diagnostics
Algorithm 1 Graph-informed variational training with localized discriminators
@@ 6.2 Synthetic Data
We consider two synthetic case studies for continuous Bayesian networks and assess the success of the graph-informed GANs using domain-grounded diagnostics. The first synthetic experiment uses the structured linear-Gaussian Hasse network from Figure 1 to isolate the effect of discriminator localization. In addition to the objective-level behavior discussed alongside Theorem 5.4, we use this example to assess whether graph-informed training preserves global distributional quality using the energy distance and how the objective functional dynamics depend on the Lipschitz constant. The second synthetic example is the ball-throwing time series, whose explicit conditional structure makes it useful for studying practical training dynamics and stability in a continuous setting.
@@ 6.2.1 Structured linear-Gaussian Hasse network
In Section 5, we assessed the structural fidelity through recovery of the parent coefficients $\{\phi_{ji}\}$ by regressing generated samples onto the true parent set and averaging the resulting $L_{2}$ error over non-root nodes. Here we complement that local diagnostic with a global one, namely the energy distance on the full eight-dimensional validation sample in Figure 3(a). Taken together with Figure 2(b), the full picture is that the graph-informed GAN improves parent-coefficient recovery while remaining competitive with the monolithic baseline in global energy distance. The misaligned multi-discriminator achieves energy-distance values of a similar order, but fails to recover the correct conditional structure. This shows that the benefit is not simply the use of several discriminators, but the use of discriminators aligned with the correct families.
@@
In summary, this case study supports the interpretation of Theorem 5.4 in a controlled setting. The graph-informed GAN performs best on recovering the correct conditional mechanisms, while still maintaining competitive global distributional quality as compared to the graph-agnostic GAN. Moreover, the Lipschitz constant study shows that varying $L$ alters both the objective’s magnitude and its training dynamics, suggesting that $L$ can serve as a practical tuning parameter to balance regularization, stability, and sensitivity in graph-informed adversarial training.
@@
Figure 3: Additional diagnostics for the structured linear-Gaussian Hasse network, complementing the objective-level comparison in Figure 2. (a) Energy distance (lower is better) provides a global measure of distributional discrepancy across the three discriminator constructs under the Lipschitz-restricted interpolative divergence (cf. parent-coefficient recovery in Figure 2(b)); boxplots summarize $13$ paired runs per model. (b) Variational loss for the graph-informed GAN trained using an interpolative divergence at several Lipschitz constants $L$, with the graph-agnostic GAN at $L=1$ for reference; mean with a min-max envelope based on 33 runs per model.

### 02 Apr 2026 v2

@@ 5.1 Infimal subadditivity for Bayesian networks over a DAG
$\begin{split}I_{1,2}[\phi](x_{1},x_{2})&+I_{1,2}[\psi](x_{1}^{\prime},x_{2}^{\prime})=\int_{\mathcal{X}_{3}}\phi(x_{1},x_{2},x_{3})Q_{3\mid 2}(dx_{3})+\int_{\mathcal{X}_{3}}\psi(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime})Q_{3\mid 2}(dx_{3}^{\prime})\\ &=\int_{\mathcal{X}_{3}}\int_{\mathcal{X}_{3}}(\phi(x_{1},x_{2},x_{3})+\psi(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}))Q(dx_{3}|x_{2})Q(dx^{\prime}_{3}|x^{\prime}_{2})\\ &\leq\int_{\mathcal{X}_{3}}\int_{\mathcal{X}_{3}}c(\bm{x},\bm{x}^{\prime})Q(dx_{3}|x_{2})Q(dx^{\prime}_{3}|x^{\prime}_{2})\end{split}$
Then, the result follows as in proof of Theorem 5.3. ∎
From the perspective of adversarial learning, the preceding theorems justify replacing a graph-agnostic GAN with a monolithic discriminator by a GiGAN with family-level discriminators. This is not only a computational simplification; since GAN training amounts to a learned two-sample testing problem (Goodfellow et al., 2014; Nowozin et al., 2016), localization decomposes a high-dimensional discrimination task into a collection of lower-dimensional subproblems that are easier to optimize and diagnose. Theorems 5.3, 5.5, and 5.6 show that these local games define principled surrogates for the corresponding graph-constrained global objectives.
@@
The same lower and upper bounds hold for IPMs. This follows directly from the monotonicity property $\mathfrak{D}_{f}^{\Gamma_{i\cup\mathrm{Pa}(i)}}(Q_{i\cup\mathrm{Pa}(i)}\|P_{i\cup\mathrm{Pa}(i)})\leq\mathfrak{D}_{f}^{\Gamma}(Q\|P)$ for all $i\in V$. The proof of the upper bound is already contained in the proof of Theorem 5.3.
@@ 5.2 Examples of admissible sets satisfying condition (C)
Let $G$ be an arbitrary DAG. Then, $f$-divergences satisfy $\frac{1}{n}$-linear generalized infimal additivity with $\mathcal{M}=\mathcal{P}(\mathcal{X})$ or $\mathcal{P}^{G}$
@@
for $\bm{x}=(x_{1},\dots,x_{n})$ and $\bm{x}^{\prime}=(x^{\prime}_{1},\dots,x^{\prime}_{n})$. Let $\Gamma^{L}=\textrm{Lip}^{L}_{b}(\mathcal{X})$. In this case, condition (C) can be satisfied if $Q_{V^{(i-1)}\mid\mathrm{Pa}(i)}\otimes\prod_{j=i+1}^{n}Q_{j\mid\mathrm{Pa}(j)}$ viewed as a function of $\bm{x}_{i\,\cup\,\mathrm{Pa}(i)}$, is Lipschitz-continuous with respect to 1-Wasserstein distance (Bartl and Eckstein, 2025; Piening and Chung, 2024; Altekrüger et al., 2023; Bénézet et al., 2024). We prove this result in Proposition B.2.
In generative modeling, the assumption that $Q_{V^{(i-1)}\mid\mathrm{Pa}(i)}\otimes\prod_{j=i+1}^{n}Q_{j\mid\mathrm{Pa}(j)}$ viewed as a function of $\bm{x}_{i\,\cup\,\mathrm{Pa}(i)}$, is Lipschitz-continuous with respect to 1-Wasserstein distance (also given in equation 42), can be interpreted as follows: Let $(\mathcal{Z},\mathcal{F},\nu)$ be a probability space, $\mathcal{X}_{\mathrm{Pa}(i)}\subset\mathbb{R}^{m}$ and $g:\mathcal{X}_{\mathrm{Pa}(i)}\times\mathcal{Z}\to\mathbb{R}^{d}$. We assume that for every $\bm{x}_{\mathrm{Pa}(i)}\in\mathcal{X}_{\mathrm{Pa}(i)}$, the map $z\mapsto g(\bm{x}_{\mathrm{Pa}(i)},z)$ is measurable and $g(\bm{x}_{\mathrm{Pa}(i)},\cdot)_{\#}\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$, and for any $\bm{x}_{\mathrm{Pa}(i)},\bm{x}^{\prime}_{\mathrm{Pa}(i)}\in\mathcal{X}_{\mathrm{Pa}(i)}$ and all $z\in\mathcal{Z}$, $\|g(\bm{x}_{\mathrm{Pa}(i)},z)-g(\bm{x}^{\prime}_{\mathrm{Pa}(i)},z)\|\leq Cd_{\mathcal{X}_{\mathrm{Pa}(i)}}(\bm{x}_{\mathrm{Pa}(i)},\bm{x}^{\prime}_{\mathrm{Pa}(i)})$. Then, a measurable generator with finite first moment and uniform Lipschitz dependence on the conditioning random vector induces a $W_{1}$-Lipschitz family of conditional distributions, see Arjovsky et al. (2017); Papamakarios et al. (2021). This is proved in Appendix B.
Therefore, such an assumption (see equation 42) expresses a quantitative stability property of the model: small perturbations of the conditioning variables induce proportionally small changes in the corresponding conditional distributions in $1$-Wasserstein distance. This stability naturally arises when conditional distributions are represented as the pushforwards of a reference noise distribution through generators that are Lipschitz in their conditioning variables.
@@ 6 Infimal subadditivity in action
Section 5 shows that when the target distribution factorizes according to a known Bayesian network over a graph $G$, the corresponding global objective function can be related to a sum of lower-dimensional objective functionals defined on graph-informed families. In adversarial learning terms, this suggests replacing a graph-agnostic GAN with a monolithic discriminator by a GiGAN with localized discriminators acting on families of variables. The aim of this section is to show how the graph-informed strategy behaves in practice across a range of structured target distributions. In particular, we find that GiGANs often improve optimization stability and the recovery of the underlying conditional structure, while also revealing trade-offs between family-level and global distributional discrepancy, and between predictive accuracy and computational cost.
@@ 22   Record validation and runtime diagnostics
Algorithm 1 GiGAN training with localized discriminators
@@ 6.2 Synthetic Data
We consider two synthetic case studies for continuous Bayesian networks and assess the success of the GiGANs using domain-grounded diagnostics. The first synthetic experiment uses the structured linear-Gaussian Hasse network from Figure 1 to isolate the effect of discriminator localization. In addition to the objective-level behavior discussed alongside Theorem 5.3, we use this example to assess whether graph-informed training preserves global distributional quality using the energy distance and how the objective functional dynamics depend on the Lipschitz constant. The second synthetic example is the ball-throwing time series, whose explicit conditional structure makes it useful for studying practical training dynamics and stability in a continuous setting.
@@ 6.2.1 Structured linear-Gaussian Hasse network
In Section 5, we assessed the structural fidelity through recovery of the parent coefficients $\{\phi_{ji}\}$ by regressing generated samples onto the true parent set and averaging the resulting $L_{2}$ error over non-root nodes. Here we complement that local diagnostic with a global one, namely the energy distance on the full eight-dimensional validation sample in Figure 3(a). Taken together with Figure 2(b), the full picture is that the GiGAN improves parent-coefficient recovery while remaining competitive with the monolithic baseline in global energy distance. The misaligned multi-discriminator achieves energy-distance values of a similar order, but fails to recover the correct conditional structure. This shows that the benefit is not simply the use of several discriminators, but the use of discriminators aligned with the correct families.
@@
In summary, this case study supports the interpretation of Theorem 5.3 in a controlled setting. The GiGAN performs best on recovering the correct conditional mechanisms, while still maintaining competitive global distributional quality as compared to the graph-agnostic GAN. Moreover, the Lipschitz constant study shows that varying $L$ alters both the objective’s magnitude and its training dynamics, suggesting that $L$ can serve as a practical tuning parameter to balance regularization, stability, and sensitivity in graph-informed adversarial training.
@@
Figure 3: Additional diagnostics for the structured linear-Gaussian Hasse network, complementing the objective-level comparison in Figure 2. (a) Energy distance (lower is better) provides a global measure of distributional discrepancy across the three discriminator constructs under the Lipschitz-restricted interpolative divergence (cf. parent-coefficient recovery in Figure 2(b)); boxplots summarize $13$ paired runs per model. (b) Variational loss for the GiGAN trained using an interpolative divergence at several Lipschitz constants $L$, with the graph-agnostic GAN at $L=1$ for reference; mean with a min-max envelope based on 33 runs per model.